# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ B.1 Noether charge
$=$ $\frac{1}{2}p^{\alpha}p^{\beta}\left(\xi^{\mu}\partial_{\mu}\bar{h}_{\alpha\beta}+\bar{h}_{\alpha\mu}\partial_{\beta}\xi^{\mu}+\bar{h}_{\beta\mu}\partial_{\alpha}\xi^{\mu}+\tau_{\alpha}\partial_{\beta}\Lambda+\tau_{\beta}\partial_{\alpha}\Lambda\right)$
$-\frac{1}{2m}\bar{h}_{\alpha\beta}p^{\alpha}p^{\beta}p^{\gamma}\mathcal{L}_{\xi}\tau_{\gamma}$
@@
To find the formula for $\delta_{\chi}\bar{h}_{\alpha\beta}$ we have used Eqs. (10)-(13). Now, since $\bar{h}_{\alpha\beta}$ is Milne-invariant, setting $v^{\mu}=u^{\mu}$ we obtain the equality
@@
Taking together Eqs. (174), (175) and (179) gives
@@
B.3 Hydrostatic conditions
Let us rephrase the Killing conditions in Eq. (60) using the Milne-invariant data as defined in Eqs. (55) and (57). Firstly,
$\delta_{\chi}\tau_{\mu}$ $=\xi^{\nu}\partial_{\nu}\tau_{\mu}+\tau_{\nu}\partial_{\mu}\xi^{\nu}=\partial_{\mu}\left(\frac{1}{T}\right)-\frac{u^{\nu}}{T}H_{\mu\nu}=0~{}\implies$ (181)
@@ B.3 Hydrostatic conditions
$\delta_{\chi}A_{\mu}$ $=\xi^{\nu}\partial_{\nu}A_{\mu}+A_{\nu}\partial_{\mu}\xi^{\nu}+\partial_{\mu}\Lambda=\partial_{\mu}\left(\frac{\mu}{mT}\right)-\frac{u^{\nu}}{T}f_{\mu\nu}=0\,,\implies$ (183)
@@
B.4 Covariant derivative of $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{n}}$
For the purposes of this Appendix we define
$\sigma_{d}=\sigma_{d+1}\delta(\tau_{\rho}p^{\rho}-m)\,.$ (187)
¯$\begin{split}\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=&\frac{1}{m^{n}}\bar{\nabla}_{\mu}\int p^{\mu}p^{\alpha_{2}}\cdots p^{\alpha_{n}}f\sigma_{d}\\ =&\bar{\Gamma}^{\mu}_{\lambda\mu}\mathcal{I}^{\lambda\alpha_{2}\cdots\alpha_{n}}+\cdots+\bar{\Gamma}^{\alpha_{n}}_{\lambda\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\lambda}+\frac{1}{m^{n}}\int p^{\mu}p^{\alpha_{2}}\cdots p^{\alpha_{n}}\bar{\nabla}_{\mu}\left(f\sigma_{d}\right)\,.\end{split}$ (188)
Now,
¯$\bar{\nabla}_{\mu}(f\sigma_{d})$ $=\partial_{\mu}f\sigma_{d}+f\bar{\nabla}_{\mu}\sigma_{d}$ (189)
Plugging it into Eq. (188), integrating the last term by parts and comparing with Eq. (29) shows
¯$\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}\left(X_{L}[f]+\frac{1}{2m^{2}}p^{2}H^{\mu}\,_{\rho}p^{\rho}\frac{\partial}{\partial p^{\mu}}f\right)\sigma_{d}\,.$ (193)
Integrating the last term by parts gives
¯$\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}X_{L}[f]\sigma_{d}-\frac{n-1}{2}H^{(\alpha_{2}}\,_{\rho}h_{\mu\nu}\mathcal{I}^{\mu\nu\rho|\alpha_{3}\cdots\alpha_{n})}-H_{\mu\nu}v^{\nu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}\,.$ (194)
Setting $v^{\mu}=u^{\mu}$ and $h_{\mu\nu}=g_{\mu\nu}$ gives Eq. (72).
@@ B.5 Conservation of energy
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}=\frac{m}{2}\mathcal{I}^{\gamma\delta\mu}\bar{\nabla}_{\mu}g_{\gamma\delta}-\frac{m}{2}g_{\mu\nu}H^{(\mu}\,_{\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\nu)}\,.$ (195)
@@
¯$\begin{split}\bar{\nabla}_{\mu}g_{\gamma\delta}&=\left(\bar{P}^{\alpha}_{\gamma}+\tau_{\gamma}u^{\alpha}\right)\bar{\nabla}_{\mu}g_{\alpha\delta}=g_{\gamma\beta}\bar{\nabla}_{\mu}\left(h^{\alpha\beta}g_{\alpha\delta}\right)-\tau_{\gamma}g_{\alpha\delta}\bar{\nabla}_{\mu}u^{\alpha}\\ &=g_{\gamma\beta}\bar{\nabla}_{\mu}\left(\delta^{\beta}_{\delta}-\tau_{\delta}u^{\beta}\right)-\tau_{\gamma}g_{\alpha\delta}\bar{\nabla}_{\mu}u^{\alpha}=-2\tau_{(\gamma}g_{\alpha\delta)}\bar{\nabla}_{\mu}u^{\alpha}\,\end{split}$ (196)
@@
Altogether we obtain
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}=-H_{\mu\rho}u^{\rho}\mathcal{E}^{\mu}-\mathcal{T}^{\mu\nu}g_{\rho(\mu}\bar{\nabla}_{\nu)}u^{\rho}\,.$ (198)
@@ B.6 Derivation of the equations of motion (129)
In this subsection, we derive the equations of motion in Eq. (129) from the evolution of the $k$-th moment of the distribution function. Shifting the projection operator inside the covariant derivative in Eq. (128) we get
@@
£$\begin{split}\pounds_{u}I^{\alpha_{1}\cdots\alpha_{k}}&=u^{\mu}\partial_{\mu}I^{\alpha_{1}\cdots\alpha_{k}}-I^{\mu\alpha_{2}\cdots\alpha_{k}}\partial_{\mu}u^{\alpha_{1}}-\cdots-I^{\alpha_{1}\cdots\alpha_{k-1}\mu}\partial_{\mu}u^{\alpha_{k}}\\ &=u^{\mu}\nabla_{\mu}I^{\alpha_{1}\cdots\alpha_{k}}-I^{\mu\alpha_{2}\cdots\alpha_{k}}\nabla_{\mu}u^{\alpha_{1}}-\cdots-I^{\alpha_{1}\cdots\alpha_{k-1}\mu}\nabla_{\mu}u^{\alpha_{k}}\,,\end{split}$ (204)
@@
(£$\left(\pounds_{u}+\nabla_{\mu}u^{\mu}\right)I^{\alpha_{1}\cdots\alpha_{k}}$ $+\nabla_{\mu}I^{\mu\alpha_{1}\cdots\alpha_{k}}+\big{[}h^{\alpha_{k}\sigma}\left(u^{\mu}I^{\alpha_{1}\cdots\alpha_{k-1}}+2I^{\mu\alpha_{1}\cdots\alpha_{k-1}}\right)+\cdots$
@@ B.7 Evaluating $S^{\alpha\beta}$ and $S^{\alpha\beta\gamma}$
In the definition of the hydrodynamic distribution function we replace the chemical potential $\mu$ with the number density $n$ with the use of Eq. (93):
@@
$\begin{split}X_{L}[f_{0}]=&\frac{p^{\mu}}{m}\partial_{\mu}f_{0}-\frac{1}{m}\Gamma^{\mu}_{\gamma\delta}p^{\gamma}p^{\delta}\frac{\partial f_{0}}{\partial p^{\mu}}\\ =&-\frac{p^{\mu}}{m}\left[\frac{1}{2mT}p^{\gamma}p^{\delta}\partial_{\mu}g_{\gamma\delta}-\frac{1}{2mT^{2}}\left(g_{\gamma\delta}p^{\gamma}p^{\delta}-dmT\right)\partial_{\mu}T-\frac{\partial_{\mu}n}{n}\right]f_{0}\\ &+\frac{1}{mT}\Gamma^{\mu}_{\gamma\delta}p^{\gamma}p^{\delta}g_{\mu\rho}p^{\rho}f_{0}\\ =&-\frac{p^{\mu}}{m}\left[\frac{1}{2mT}p^{\gamma}p^{\delta}\nabla_{\mu}g_{\gamma\delta}-\frac{1}{2mT^{2}}\left(g_{\gamma\delta}p^{\gamma}p^{\delta}-dmT\right)\partial_{\mu}T-\frac{\partial_{\mu}n}{n}\right]f_{0}\,.\end{split}$ (207)
@@
Using the equations of motion in Eq. (102) we can eliminate the terms containing $u^{\mu}\partial_{\mu}n$, $u^{\mu}\nabla_{\mu}u^{\delta}$, and $u^{\mu}\partial_{\mu}T$. This way we obtain
@@
where we also use the change of variables $p^{\mu}\rightarrow(p^{\mu}+mu^{\mu})$ and use Eq. (212). Similarly, using the zeroth and first moment we calculate the second moment as

### v2

@@ B.1 Noether charge
$=$ $\frac{1}{2}p^{\alpha}p^{\beta}\left(\xi^{\mu}\partial_{\mu}\bar{h}_{\alpha\beta}+\bar{h}_{\alpha\mu}\partial_{\beta}\xi^{\mu}+\bar{h}_{\beta\mu}\partial_{\alpha}\xi^{\mu}+\tau_{\alpha}\partial_{\beta}\Lambda+\tau_{\beta}\partial_{\alpha}\Lambda\right)-\frac{1}{2m}\bar{h}_{\alpha\beta}p^{\alpha}p^{\beta}p^{\gamma}\mathcal{L}_{\xi}\tau_{\gamma}$
@@
To find the formula for $\delta_{\chi}\bar{h}_{\alpha\beta}$, we have used Eqs. (11)-(14). Now, since $\bar{h}_{\alpha\beta}$ is Milne-invariant, setting $v^{\mu}=u^{\mu}$ we obtain the equality
@@
Taking together Eqs. (176), (177), and (181), gives
@@
B.3 Covariant derivative of $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{n}}$
For the purposes of this Appendix, we define
$\sigma_{d}=\sigma_{d+1}\delta(\tau_{\rho}p^{\rho}-m)\,.$ (183)
$\begin{split}\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=&\frac{1}{m^{n}}\bar{\nabla}_{\mu}\int p^{\mu}p^{\alpha_{2}}\cdots p^{\alpha_{n}}f\sigma_{d}\\ =&\bar{\Gamma}^{\mu}_{\lambda\mu}\mathcal{I}^{\lambda\alpha_{2}\cdots\alpha_{n}}+\cdots+\bar{\Gamma}^{\alpha_{n}}_{\lambda\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\lambda}+\frac{1}{m^{n}}\int p^{\mu}p^{\alpha_{2}}\cdots p^{\alpha_{n}}\bar{\nabla}_{\mu}\left(f\sigma_{d}\right)\,.\end{split}$ (184)
Now,
$\bar{\nabla}_{\mu}(f\sigma_{d})$ $=\partial_{\mu}f\sigma_{d}+f\bar{\nabla}_{\mu}\sigma_{d}$ (185)
Plugging it into Eq. (184), integrating the last term by parts, and comparing with Eq. (44), shows that
$\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}\left(X_{L}[f]+\frac{1}{2m^{2}}p^{2}H^{\mu}\,_{\rho}p^{\rho}\frac{\partial}{\partial p^{\mu}}f\right)\sigma_{d}\,.$ (189)
Integrating the last term by parts gives
$\bar{\nabla}_{\mu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}X_{L}[f]\sigma_{d}-\frac{n-1}{2}H^{(\alpha_{2}}\,_{\rho}h_{\mu\nu}\mathcal{I}^{\mu\nu\rho|\alpha_{3}\cdots\alpha_{n})}-H_{\mu\nu}v^{\nu}\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}\,.$ (190)
Setting $v^{\mu}=u^{\mu}$ and $h_{\mu\nu}=g_{\mu\nu}$ gives Eq. (64).
B.4 Hydrostatic conditions
Let us rephrase the Killing conditions in Eq. (82) using the Milne-invariant data defined in Eqs. (77) and (79). Firstly,
$\delta_{\chi}\tau_{\mu}$ $=\xi^{\nu}\partial_{\nu}\tau_{\mu}+\tau_{\nu}\partial_{\mu}\xi^{\nu}=\partial_{\mu}\left(\frac{1}{T}\right)-\frac{u^{\nu}}{T}H_{\mu\nu}=0$$\implies$ (191)
@@ B.4 Hydrostatic conditions
$\delta_{\chi}A_{\mu}$ $=\xi^{\nu}\partial_{\nu}A_{\mu}+A_{\nu}\partial_{\mu}\xi^{\nu}+\partial_{\mu}\Lambda=\partial_{\mu}\left(\frac{\mu}{mT}\right)-\frac{u^{\nu}}{T}f_{\mu\nu}=0\,$$\implies$ (193)
@@ B.5 Conservation of energy
$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}=\frac{m}{2}\mathcal{I}^{\gamma\delta\mu}\bar{\nabla}_{\mu}g_{\gamma\delta}-\frac{m}{2}g_{\mu\nu}H^{(\mu}\,_{\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\nu)}\,.$ (197)
@@
$\begin{split}\bar{\nabla}_{\mu}g_{\gamma\delta}&=\left(\bar{P}^{\alpha}_{\gamma}+\tau_{\gamma}u^{\alpha}\right)\bar{\nabla}_{\mu}g_{\alpha\delta}=g_{\gamma\beta}\bar{\nabla}_{\mu}\left(h^{\alpha\beta}g_{\alpha\delta}\right)-\tau_{\gamma}g_{\alpha\delta}\bar{\nabla}_{\mu}u^{\alpha}\\ &=g_{\gamma\beta}\bar{\nabla}_{\mu}\left(\delta^{\beta}_{\delta}-\tau_{\delta}u^{\beta}\right)-\tau_{\gamma}g_{\alpha\delta}\bar{\nabla}_{\mu}u^{\alpha}=-2\tau_{(\gamma}g_{\alpha\delta)}\bar{\nabla}_{\mu}u^{\alpha}\,\end{split}$ (198)
@@
Altogether, we obtain
$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}=-H_{\mu\rho}u^{\rho}\mathcal{E}^{\mu}-\mathcal{T}^{\mu\nu}g_{\rho(\mu}\bar{\nabla}_{\nu)}u^{\rho}\,.$ (200)
@@ B.6 Derivation of the equations of motion (132)
In this subsection, we derive the equations of motion in Eq. (132) from the evolution of the $k$-th moment of the distribution function. Shifting the projection operator inside the covariant derivative in Eq. (131), we get
@@
[…]
@@
$\left(\pounds_{u}+\nabla_{\mu}u^{\mu}\right)I^{\alpha_{1}\cdots\alpha_{k}}$ $+\nabla_{\mu}I^{\mu\alpha_{1}\cdots\alpha_{k}}+\big[h^{\alpha_{k}\sigma}\left(u^{\mu}I^{\alpha_{1}\cdots\alpha_{k-1}}+2I^{\mu\alpha_{1}\cdots\alpha_{k-1}}\right)+\cdots$
@@ B.7 Evaluating $S^{\alpha\beta}$ and $S^{\alpha\beta\gamma}$
In the definition of the hydrodynamic distribution function $f_{0}$, we replace the chemical potential $\mu$ with the number density $n$ with the use of Eq. (94):
@@
$\begin{split}X_{L}[f_{0}]=&\frac{p^{\mu}}{m}\partial_{\mu}f_{0}-\frac{1}{m}\Gamma^{\mu}_{\gamma\delta}p^{\gamma}p^{\delta}\frac{\partial f_{0}}{\partial p^{\mu}}\\ =&-\frac{p^{\mu}}{m}\left[\frac{1}{2mT}p^{\gamma}p^{\delta}\partial_{\mu}g_{\gamma\delta}-\frac{1}{2mT^{2}}\left(g_{\gamma\delta}p^{\gamma}p^{\delta}-dmT\right)\partial_{\mu}T-\frac{\partial_{\mu}n}{n}\right]f_{0}+\frac{1}{mT}\Gamma^{\mu}_{\gamma\delta}p^{\gamma}p^{\delta}g_{\mu\rho}p^{\rho}f_{0}\\ =&-\frac{p^{\mu}}{m}\left[\frac{1}{2mT}p^{\gamma}p^{\delta}\nabla_{\mu}g_{\gamma\delta}-\frac{1}{2mT^{2}}\left(g_{\gamma\delta}p^{\gamma}p^{\delta}-dmT\right)\partial_{\mu}T-\frac{\partial_{\mu}n}{n}\right]f_{0}\,.\end{split}$ (209)
@@
Using the equations of motion in Eqs. (103)-(105), we can eliminate the terms containing $u^{\mu}\partial_{\mu}n$, $u^{\mu}\nabla_{\mu}u^{\delta}$, and $u^{\mu}\partial_{\mu}T$. This way we obtain
@@
where we also use the change of variables $p^{\mu}\rightarrow(p^{\mu}+mu^{\mu})$ and use Eq. (214). Similarly, using the zeroth and first moment, we calculate the second moment as